# Neural Graph Matching Improves Retrieval Augmented Generation in Molecular Machine Learning

### v1

@@ 5 Conclusion
Software and Data
The code that reproduces results in this paper will be made publicly available upon peer review.
The NIST (2020) database is a commercial database widely used by domain experts. This database is not freely available, and only purchasable through several vendors worldwide [^3] https://chemdata.nist.gov/dokuwiki/doku.php?id=chemdata:distributors. Given the scale of effort required to purchase samples, run experiments, and collect the 530K spectra, and that NIST (2020) is the only database where all spectra have collision energy annotations, this dataset is a reasonable investment in mass spectrum-related research in the absence of a thorough open-source replacement. This commercial dataset is also used in many mass spectrum methodology papers and peer methods, including but not limited to Goldman et al. (2023, 2024); Li et al. (2021); Murphy et al. (2023); Young et al. (2024a, b). Given the license-based access of the dataset, our pretrained model weights will be made available to research groups with a purchased NIST license.
@@ Appendix B Retrieval Accuracy with [M+H]+ Only
One design choice in MS simulator development is how many adduct types are supported by the model. Our MARASON aims to cover most adduct types in the NIST (2020) database, while some methods e.g. FraGNNet (Young et al., 2024b) only support the most common adduct type—[M+H]+. It is still feasible to compare these methods by restricting testing adduct type as [M+H]+, where the retrieval accuracy is shown as follows in Table 4. The performance of FraGNNet is quoted from their paper (therefore missing accuracies at $k=2,4,8$, also missing error bars), and other baselines are implemented with our benchmark. It is worth noting that although Young et al. (2024b) report better retrieval accuracy of FraGNNet than all peer methods, our reevaluation of all methods on the [M+H]+-only subset shows that ICEBERG (Goldman et al., 2024) still outperforms, suggesting that single-bond breaking in MS simulator design might not be superior to multiple-bond breaking. Incorporating collision energy for ICEBERG also improves the retrieval accuracy on [M+H]+, and our MARASON reaches state-of-the-art retrieval accuracy.
@@
ICEBERG (w/ collision energy) 0.270$\pm$0.016 0.487±$\pm 0.009$ 0.611$\pm$0.011 0.679$\pm$0.013 0.735$\pm$0.013 0.840$\pm$0.009 0.877$\pm$0.008
@@ Appendix D Evaluation of Matching Patterns: Visualizing Failure Modes of Traditional Graph Matching
In Fig. 5, the top 50 matched pairs, which exhibit higher similarity scores, predominantly correspond to exact structural matches. In contrast, the lower-scored pairs include more fragments that, while structurally distinct, follow similar fragmentation pathways. Notably, the learned matching function in MARASON effectively identifies an informative and chemically interpretable fragment-matching strategy. This ensures that the retrieved intensity information is accurately assigned to the correct target fragments, thereby enhancing MS simulation accuracy.

### v2

@@ 4.2.5 Qualitative Study of Matching Patterns
4.2.6 What if there is no reference structure?
To understand the relationship between RAG’s improvement and the availability of a similar-enough reference structure, for all testing data in NIST (2020), we group structural pairs based on their Tanimoto similarities and compare performances of MARASON with or without RAG (same model configurations as in Table 3). As shown in Table 4, RAG starts to bring a significant performance gain when the Tanimoto similarity is larger than 0.3. When the best reference structure has a Tanimoto similarity between 0.1 and 0.3, RAG does not bring a significant improvement, while MARASON still performs robustly with such irrelevant references. The slight performance drop between Tanimoto similarity (0.1, 0.2] also suggests a simple trick to further improve MARASON’s accuracy: use standard MARASON when the retrieved structure has a Tanimoto similarity $>0.2$ and use the non-RAG version otherwise.
@@ Acknowledgements
Impact Statement
This paper presents work that aims to advance both machine learning and mass spectroscopy. As a methodology-focused paper, this paper itself does not pose any significant societal impact that is required to be highlighted here. The authors are also aware that mass spectroscopy could be used in human-centric research campaigns such as metabolomics, and it is important to ensure that the reference database is not biased towards ethnicity, sex, and age, or at least being acknowledged about the potential bias. On the other hand, efforts to develop better mass spectroscopy models, including this paper, will potentially lower the cost and expedite those discovery campaigns. If such models are used properly for the underrepresented cohort, it will finally mitigate the bias and unfairness in human-centric studies.
@@ Appendix B Retrieval Accuracy with [M+H]+ Only
One design choice in MS/MS simulator development is how many adduct types are supported by the model. Our MARASON aims to cover most adduct types in the NIST (2020) database, while some methods e.g. FraGNNet (Young et al., 2024b) only support the most common adduct type—[M+H]+. It is still feasible to compare these methods by restricting testing adduct type as [M+H]+, where the retrieval accuracy is shown as follows in Table 6. The performance of FraGNNet is quoted from their paper (therefore missing accuracies at $k=2,4,8$, also missing error bars), and other baselines are implemented with our benchmark. It is worth noting that although Young et al. (2024b) report better retrieval accuracy of FraGNNet than all peer methods, our reevaluation of all methods on the [M+H]+-only subset shows that ICEBERG (Goldman et al., 2024) still outperforms, suggesting that single-bond breaking in MS/MS simulator design might not be superior to multiple-bond breaking. Incorporating collision energy for ICEBERG also improves the retrieval accuracy on [M+H]+, and our MARASON reaches state-of-the-art retrieval accuracy.
@@
ICEBERG (w/ collision energy) 0.270$\pm$0.016 0.487$\pm$0.009 0.611$\pm$0.011 0.679$\pm$0.013 0.735$\pm$0.013 0.840$\pm$0.009 0.877$\pm$0.008
@@ Appendix D Evaluation of Matching Patterns: Visualizing Failure Modes of Traditional Graph Matching
In Fig. 6, the top 50 matched pairs, which exhibit higher similarity scores, predominantly correspond to exact structural matches. In contrast, the lower-scored pairs include more fragments that, while structurally distinct, follow similar fragmentation pathways. Notably, the learned matching function in MARASON effectively identifies an informative and chemically interpretable fragment-matching strategy. This ensures that the retrieved intensity information is accurately assigned to the correct target fragments, thereby enhancing MS/MS simulation accuracy.